d^y/dx^-4y=xsinhx

1
Expert's answer
2021-11-19T00:36:21-0500

Solution;

"\frac{d^2y}{dx^2}-4y=xsinhx"

For the complementary solution:

The characteristic polynomial is;

"m^2-4=0"

"m^2=4"

"m=\displaystyle _-^+2"

Hence the homogeneous solution is;

"y_h=c_1e^{2x}+c_2e^{-2x}"

For the particular Integral;

"P.I=\frac{1}{D^2-4}xsinhx"

"=\frac{1}{(D+2)(D-2)}x\frac{e^x-e^{-x}}{2}"

"=\frac{e^x}{2(D+2)(D-2)}x-\frac{e^{-x}}{2(D+2)(D-2)}"

"=\frac{e^x}{2(D+1+2)(D+1-2)}x-\frac{e^{-x}}{2(D-1+2)(D-1-2)}x"

"=\frac{e^x}{2(D^2+2D-3)}x-\frac{e^{-x}}{2(D^2-2D-3)}x"

"=\frac{e^x}{2}[\frac{1}{-3(1-\frac{D^2+2D}{3})}]x-\frac{e^{-x}}{2}[\frac{1}{-3(1-\frac{(2D-D^2)}{3})}]x"

"=\frac{e^x}{-6}[1+\frac{D^2+2D}{3}+...]x+\frac{e^{-x}}{6}[1-\frac{2D-D^2}{3}+...]x"

"=\frac{-e^x}{6}[x+\frac23]+\frac{e^{-x}}{6}[x+\frac23]"

"=\frac{-x}{3}(\frac{e^x-e^{-x}}{2})-\frac29(\frac{e^x-e^{-x}}{2})"

Hence;

"P.I=\frac{-x}{3}sinhx-\frac29sinhx" ="-\frac19(3x+2)sinhx"

Complete solution is ;

"y=C.F+P.I"

"y=c_1e^{2x}+c_2e^{-2x}-\frac 19(3x+2)sinhx"